OpenStudy (anonymous):

simplify: 4 divided by the square root of 24x

OpenStudy (anonymous):

\[4\div \sqrt{24x} = 1 \] divide both sides by 4 \[\sqrt{24x} = \frac{ 1 }{ 4 }\] square both sides \[24x = \frac{ 1 }{ 16 }\] divide both sides by 24 \[x = \frac{ 1 }{ 384 }\] You can't really do anything unless it is equal to something so I put in a 1 If it is not equal to anything the simplification would be\[\sqrt{\frac{ 2 }{ 3 }}\div \sqrt{x}\]